If 8 men can work in 12 days to assemble 16 machines, how long will it take 15 men to assemble 60 machines?
Answer by richwmiller(17219)

3*12=96 man days for 16 machines
96/16=6 man days per machine
60 *6=360 man days for 60 machines
360/15=24 days for 15 men to assemble 60 machines